OpenStudy (anonymous):

Solve the quadratic equation. (x + 2)2 = 36 A. x = ±6 B. x = 4 and x = -8 C. x = ±√6 D. x = ±√2

OpenStudy (anonymous):

first you square root both sides to get \[x+2=\pm6\] then put x+2 equal to both -6 and +6 \[x+2=-6\] \[x+2=6\] in both you subtract two and get x=-8 and x=4